Circular No. 8627

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2005 V1 (BERNARDI) D. Tholen, University of Hawaii, reports the discovery of a comet by

Fabrizio Bernardi on images taken with the Canada-France-Hawaii Telescope (discovery observation below; the observing was assisted by A. Boattini, T. Burdullis, and M. Laychak, with Tholen as principal investigator of the University of Hawaii Asteroid Search project). The comet shows a very faint tail extending as much as 12′′ in p.a. ∼ 290^{o} on the discovery images, and the magnitude was made in a 3′′.7 photometric aperture. Images obtained by H. Hsieh on Nov. 4.6 UT with the 2.24-m telescope at Mauna Kea also show the tail. The point-spread function of the comet’s head was consistently ∼ 0′′.3 larger than stars of similar brightness.

2005 UT  α_{2000}  δ_{2000}  Mag.
Nov. 1.62042  11^{h}33^{m}44.^{s}69  +8^{o}23′02′′.0  20.5

The available astrometry, the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2005-V41.

T = 2005 Sept. 3.113 TT  ω = 4.^{o}483
e = 0.62125  Ω = 130.300 }2000.0
q = 2.44635 AU  i = 15.574
a = 6.45896 AU n^{o} = 0.060043 P = 16.42 years

(1862) APOLLO S. J. Ostro, L. A. M. Benner, J. D. Giorgini, Jet Propulsion Laboratory;

M. C. Nolan, A. A. Hine, E. S. Howell, Arecibo Observatory; J. L. Margot, Cornell University; C. Magri, University of Maine, Farmington; and M. K. Shepard, Bloomsburg University of Pennsylvania, report that Arecibo delay-Doppler radar observations of (1862) on Oct. 30 and Nov. 1–2 show that this minor planet has a satellite whose average diameter is within a factor of two of 75 m and whose maximum orbital distance exceeds 3 km.

SUPERNOVAE 2005hk–2005in CBETs 268 and 272 contain information on 30 supernovae found by the

Sloan Digital Sky Survey II collaboration during Sept. 3–Oct. 23 (communicated by J. Frieman), all in the magnitude range g = 19–23 at discovery. All are either confirmed or suspected type-Ia supernovae except 2005hl and 2005hm, which are of type Ib. SN 2005hk (cf. IAUC 8625) was independently discovered by the SDSS II group.

2005 November 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green